x2 + y2 = 9
x2 - y2 = 1

Answer by Alan3354(69443)

x2 + y2 = 9
x2 - y2 = 1
-------------- Add
2x^2 = 10
x = ± sqrt(5)
------------
y = ± 2
-------
--> 4 points of intersection